OpenStudy (anonymous):

[2.03] Solve: 4(x + 1) = 4x + 4 No Solution 0 Any Real Number 1

OpenStudy (anonymous):

any real number satisfies the above equation hence that is the answer

OpenStudy (anonymous):

Try 0 and try 1.

OpenStudy (anonymous):

no even x=2,3,4..........satisfy the given equation

OpenStudy (anonymous):

Yes, keep giving out answers to questions without the asker learning anything.

OpenStudy (anonymous):

@supanick14 What happens if you let \(x=0\)?

OpenStudy (anonymous):

directly from my online schooling

OpenStudy (anonymous):

hope that helps.

OpenStudy (anonymous):

number one (no solution) is wrong,because every equations like this have a solution (except some of them that you don't need to know) and you know that 4(x+1) = 4x+1 so we have: 4x + 1 = 4x + 1 any idea?

OpenStudy (anonymous):

"every equations like this have a solution (except some of them that you don't need to know" Every integer is even (except some of them that you don't need to know).

OpenStudy (anonymous):

@wio is a master.

OpenStudy (anonymous):

no no @wio i don' mean that... i knew it but i think if @supanick14 think about it this way it's better :)

OpenStudy (anonymous):

I think the best way to think about it is to try out a couple points.

OpenStudy (anonymous):

If those test points work, then you can try to see if they're equivalent.

OpenStudy (anonymous):

yes,this is a nice idea too ;)

OpenStudy (wolf1728):

4x + 4 can be transformed into 4x = -4 x is easy to solve now don't you think?

OpenStudy (anonymous):

any real number

OpenStudy (anonymous):

@supanick14:May I help ?:) Got it ?:)